# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 3.3.1. Comparison with Tian’s orthogonality condition
Consider the space $\mathcal{H}(X,\omega_{0})/\mathbb{R}$ of all Kähler potentials, modulo additive constants. The group $\mathcal{G}$ naturally acts from the right on $\mathcal{H}(X,\omega_{0})/\mathbb{R},$ by demanding that $g^{*}\omega_{u}=\omega_{0}+dd^{c}(u\cdot g).$ Now assume that $\omega_{0}$ is Kähler–Einstein and consider the following orthogonality condition on $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$:
@@
where $\Delta$ denotes the Laplacian with respect to $\omega$ and $\lambda_{1}$ denotes its smallest strictly positive eigenvalue. This condition appears in Tian’s conjecture [89, Conj 5.5], settled in [48], saying that if $\omega_{0}$ is Kähler–Einstein, then there exists positive constants $\epsilon$ and $C$ such that
@@
for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.31.222Strictly speaking, in the original conjecture [89, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [48]. It should be emphasized that the functional $(\mathcal{I}-\mathcal{J})$ appearing in the right hand side above may as well be replaced by the functionals $\mathcal{I}$ or $\mathcal{J}$ (up to changing the constant $\epsilon),$ since all three functionals $\mathcal{I},(\mathcal{I}-\mathcal{J})$ and $\mathcal{J}$ are comparable (up to multiplicative constants [23, Section 2]). The inequality 3.32 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.32 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.31 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [95, Lemma A.2]:
@@
Let $X$ be a Fano manifold and assume that the Futaki character of $X$ vanishes and that the group $\mathcal{G}$ is reductive. Fix a Kähler form $\omega_{0}$ that is invariant under a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}.$ Given $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ the following two conditions are equivalent for $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}:$
$(i)\,\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g)=\mathcal{J}(v),\,\,\,\,(ii)\int_{X}v\Delta\boldsymbol{m}\omega_{0}^{n}=0,$
where $\Delta$ denotes the Laplacian with respect to the Kähler metric $\omega_{0}.$ Moreover, there exists a unique $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying any of these two conditions. When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.31 for $v.$
@@
First observe that $\mathcal{J}(u)$ is $\mathcal{K}$–invariant, since $\omega_{0}$ is assumed $\mathcal{K}$–invariant. It thus follows, precisely as in the proof of Prop 3.4, that the inf over $\mathcal{G}$ in $(i)$ may be replaced by the inf over $J\mathfrak{k}.$ The pullback of $\mathcal{J}$ to $J\mathfrak{k}$ is proper and strictly convex on $J\mathfrak{k}.$ Indeed, for any complex one-parameter curve $\phi^{\tau}$ in $\mathcal{H}(-K_{X})$ satisfying the complex geodesic equation 3.25 a direct calculation gives
$V\frac{\partial^{2}\mathcal{J}(\phi^{\tau}-\phi_{0})}{\partial\tau\partial\overline{\tau}}=\int_{X}\frac{\partial^{2}\phi^{\tau}}{\partial\tau\partial\overline{\tau}}\omega_{0}^{n}=\int_{X}\left|\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)\right|_{dd^{c}\phi^{\tau}}^{2}\omega_{0}^{n}$
and we can thus proceed precisely as in the proof of Prop 3.4. Hence, the inf in $(i)$ is attained for a unique $v\in u\cdot\mathcal{G}\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ and this $v$ is the unique critical point of the the pullback of $\mathcal{J}$ to $J\mathfrak{k}.$ Now, fix $W\in\mathfrak{k}$ and $v\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ and set $v_{t}\coloneqq v\cdot\exp(tJW).$ Then
(3.33) $V\frac{d\mathcal{J}(v_{t})}{dt}_{|t=0}=\int_{X}\left.\frac{dv_{t}}{dt}\right|_{t=0}\omega_{0}^{n}=\int_{X}\left\langle dv,JW\right\rangle\omega_{0}^{n}=\int_{X}\nabla v\cdot JW\omega_{0}^{n}$
using, in the first equality, that the Futaki character vanishes (formula 3.17). Expressing $\left\langle dv,JW\right\rangle$ as the scalar product $\nabla v\cdot JW$ between the vector field $JW$ and the gradient $\nabla v$ of $v$ with respect to the Kähler metric $\omega_{0}$. Using that $JW$ is proportional to $\nabla h_{W},$ where $h_{W}$ is the Hamiltonian function corresponding to $W,$ we deduce that
@@
Hence, $\frac{d}{dt}\mathcal{J}(v_{t})=0$ for all $W\in\mathfrak{k}$ iff condition $(ii)$ holds, as desired. Finally, when $\omega_{0}$ is Kähler–Einstein $\Delta h_{W}=\lambda_{1}h_{W}$ for any $W\in\mathfrak{k},$ where $\lambda_{1}$ denotes the smallest strictly positive eigenvalue of $\Delta$ [8, Thm 2.2]. Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.31, when $\omega_{0}$ is Kähler–Einstein. ∎
In the case when $n=1,$ so that $X=\mathbb{P}^{1}$, the functionals $\mathcal{I}-\mathcal{J}$ and $\mathcal{J}$ coincide and thus the moment condition appearing in Theorem 3.7 then coincides with condition $(ii)$ in the previous lemma.
@@ 3.4. The reduced analytic stability thresholds of K-polystable log Fano curves
Given $t<\alpha(X,\Delta)_{0}$ it will be enough to show that the functional $\phi\mapsto\mathcal{I}_{t,\mu_{0}}(\phi_{\mu})$ is bounded from above on the subspace $K$ of $\text{PSH(}X,-K_{(X,\Delta)})$ consisting of all potentials $\phi$ of $\mu\in\mathcal{P}(X)_{0}.$ To this end first note that $K$ is a compact subspace of $\text{PSH(}X,-K_{(X,\Delta)}),$ endowed with the $L^{1}$–topology. Indeed, $\mathcal{P}(X)_{0}$ is a closed compact subset of the compact space $\mathcal{P}(X),$ hence compact and, as is well-known, the map $\mu\mapsto\phi_{\mu}$ is a homeomorphism. It will thus be enough to show that $\mathcal{I}_{t,\mu_{0}}$ is continuous on $K.$ But since $\mathcal{I}_{t+\delta,\mu_{0}}$ is pointwise finite on $K$ for $\delta$ a sufficiently small positive number, this is shown precisely as in the proof of [10, Cor A.2], using the corresponding local result of Demailly–Kollar.
@@
where $F(-\gamma)_{0}$ is defined by formula 3.15. Moreover, $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ and $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}$ when $w\neq 0.$
@@
Let us first show that for any log Fano curve $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$ We will adapt the argument in [10] (concerning the case $\mathcal{G}$ trivial) to the present setup. Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.16 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
@@
To this end recall that, $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ is independent of the reference $\phi_{0}$ (by Prop 3.6, using that the Futaki character of $(X,\Delta_{w})$ vanishes). Hence, we can take $\phi_{0}$ to be induced by the Fubini–Study metric. It will be enough to show that there exists a curve $\mu^{t}$ in $\mathcal{P}(X)_{0}$ such that for any $\gamma>2,$ $F_{\gamma}(\mu^{t})\rightarrow$-$\infty$ as $t\rightarrow\infty.$ To this end first note that if $\mu$ is $S^{1}$–invariant under the standard $S^{1}$–action of $\mathbb{P}^{1}$ fixing 0 and $\infty$ and $\mathbb{Z}_{2}$–invariant under the $\mathbb{Z}_{2}$–action $(x_{1},x_{2},x_{3})\mapsto(-x_{1},-x_{2},-x_{3})$ then $\boldsymbol{m}(x_{1},x_{2},x_{3})=-\boldsymbol{m}(-x_{1},-x_{2},-x_{3})$ and it follows that $\mu\in\mathcal{P}(X)_{0}.$ We will construct $\mu^{t}$ as $\mu^{t}=dd^{c}\phi_{t}/V,$ where $\phi_{t}$ is an $S^{1}$–invariant geodesic in $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}},$ using the notation in [16] (recalled in the appendix), where such a curve $\phi_{t}$ is realized as the Legendre transform of a function $v_{t}(y)$ on the moment polytope P=[V/2,V/2]P=[-V/2,V/2] of $(X,-K_{(X,\Delta)})$ of the form
@@
where $v$ is a piecewise affine convex function on [1,1][-1,1] and $v_{0}$ is the Legendre transform of $\phi_{0}.$ It follows from well-known general results that such a curve $\phi_{t}$ is a geodesic ray in $\mathcal{H}(-K_{X})_{1,1}^{S^{1}}$ (see [41, Section 7]). Here we will take
@@
$F_{\gamma}(\mu)=\gamma\frac{\mathcal{L}(v)}{V}+(1-\gamma)\frac{V}{2}\int_{P}\left|v^{\prime}\right|dy-\int_{P}v^{\prime\prime}dy+O(1),\,\,\,\frac{\mathcal{L}(v)}{V}\coloneqq\int_{\partial P}\frac{v}{2}-\int_{P}v\frac{dy}{V},$
where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 9.1, combined with 9.3). In particular,
@@
$F_{\gamma}(\mu^{t})=\gamma t\frac{V}{4}+(1-\gamma)t\frac{V}{2}+O(1)=tV\left(\gamma\left(\frac{1}{4}-\frac{1}{2}\right)+\frac{1}{2}\right)+O(1),$
@@
The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 9.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
@@
The proof of the previous theorem shows that when $\gamma>\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ the functional $F_{\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is destabilized by a geodesic ray, i.e. $F_{\gamma}(\mu_{t})\rightarrow-\infty,$ as $t\rightarrow\infty,$ for a curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ of the form $\mu_{t}=\omega_{u_{t}}/V,$ where $u_{t}$ is a geodesic ray in $\mathcal{H}_{1,1}(X,\omega_{0}).$ This is reminiscent of the general existence/properness principle in [48, Section 3]. However, in the present setup the function $F_{\gamma}(\omega_{u_{t}}/V)$ need not be convex in $t$ (since $\gamma>1)$ which is required in [48]. Moreover, our geodesics are required to satisfy the constraint $\int_{X}\boldsymbol{m}\omega_{u_{t}}=0.$

### v3

@@ 3.3.1. Comparison with Tian’s orthogonality condition
Consider the space $\mathcal{H}(X,\omega_{0})/\mathbb{R}$ of all Kähler potentials, modulo additive constants. Assume that $\omega_{0}$ is Kähler–Einstein and consider the following orthogonality condition on $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$:
@@
where $\Delta$ denotes the Laplacian with respect to $\omega$ and $\lambda_{1}$ denotes its smallest strictly positive eigenvalue. This condition appears in Tian’s conjecture [107, Conj 5.5], settled in [61], saying that if $\omega_{0}$ is Kähler–Einstein, then there exists positive constants $\epsilon$ and $C$ such that
@@
for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.30.222Strictly speaking, in the original conjecture [107, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [61]. It should be emphasized that the functional $(\mathcal{I}-\mathcal{J})$ appearing in the right hand side above may as well be replaced by the functionals $\mathcal{I}$ or $\mathcal{J}$ (up to changing the constant $\epsilon),$ since all three functionals $\mathcal{I},(\mathcal{I}-\mathcal{J})$ and $\mathcal{J}$ are comparable (up to multiplicative constants [25, Section 2]). The inequality 3.31 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.31 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.30 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [113, Lemma A.2]:
@@
Let $X$ be a Fano manifold and assume that the Futaki character of $X$ vanishes and that the group $\mathcal{G}$ is reductive. Fix a Kähler form $\omega_{0}$ that is invariant under a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}.$ Given $u\in\mathcal{E}^{1}(X,\omega_{0})/\mathbb{R}$ the following conditions are equivalent for $v\in u\cdot G\subset\mathcal{E}^{1}(X,\omega_{0})/\mathbb{R}:$
$(i)\,\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g)=\mathcal{J}(v),\,\,\,(ii)\int_{X}v\Delta\boldsymbol{m}\omega_{0}^{n}=0,\,\,\,(iii)\int_{X}\boldsymbol{m}\omega_{v}\wedge\omega_{0}^{n-1}=0$
where $\Delta$ denotes the Laplacian with respect to the Kähler metric $\omega_{0}.$ Moreover, modulo the action of $\mathcal{K},$ there exists a unique $v\in u\cdot G\subset\mathcal{E}^{1}(X,\omega_{0})/\mathbb{R}$ satisfying any of these two conditions. When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.30 for $v.$
@@
First observe that $\mathcal{J}(u)$ is $\mathcal{K}$–invariant, since $\omega_{0}$ is assumed $\mathcal{K}$–invariant. It thus follows, precisely as in the proof of Prop 3.4, that the inf over $\mathcal{G}$ in $(i)$ may be replaced by the inf over $J\mathfrak{k},$ under the identification of $\mathcal{K}\backslash\mathcal{G}$ with $J\mathfrak{k}.$ Now fix a non-trivial real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g}.$ Let us first show that
(3.32) $V\frac{d\mathcal{J}(v\cdot\exp(-tJW))}{dt}_{|t=0}=c_{n}\int_{X}v\Delta h_{W}\omega_{0}^{n}$
for a positive constant $c_{n}$ only depending on $n.$ It is enough to consider the case when $v\in\mathcal{H}(X,-K_{X}),$ using an approximation argument, precisely as in the proof of Prop 3.4. Setting $v_{t}\coloneqq v\cdot\exp(-tJW),$
(3.33) $V\frac{d\mathcal{J}(v_{t})}{dt}_{|t=0}=\int_{X}\left.\frac{dv_{t}}{dt}\right|_{t=0}\omega_{0}^{n}=-\int_{X}\left\langle dv,JW\right\rangle\omega_{0}^{n}=-\int_{X}\nabla v\cdot JW\omega_{0}^{n}$
using, in the first equality, that the Futaki character vanishes (formula 3.19). Expressing $\left\langle dv,JW\right\rangle$ as the scalar product $\nabla v\cdot JW$ between the vector field $JW$ and the gradient $\nabla v$ of $v$ with respect to the Kähler metric $\omega_{0}$ and using that $JW=c_{n}\nabla h_{W}$ for a positive constant $c_{n}$ (where $h_{W}$ denotes the Hamiltonian function corresponding to $W)$ we deduce that
@@
In particular, $\frac{d}{dt}\mathcal{J}(v_{t})=0$ for all $W\in\mathfrak{k}$ iff condition $(ii)$ in the present lemma holds. To prove that $(i)$ is equivalent to $(ii),$ together with the uniqueness statement in the lemma, it is thus enough to show that $\mathcal{J}(v_{t})$ is strictly convex for any given $W\in\mathfrak{k}.$ To this end we first note that formula 3.32 can be reformulated as
$V\frac{d\mathcal{J}(v\cdot\exp(-tJW))}{dt}_{|t=0}=c_{n}\int_{X}h_{W}(\Delta v+n)\omega_{0}^{n-1}=\frac{c_{n}}{(n-1)!}\int_{X}h_{W}\omega_{v}\wedge\omega_{0}^{n-1},$
showing that condition $(ii)$ is also equivalent to condition $(iii).$ Indeed, integrating by parts, we can replace the integrand $v\Delta h_{W}$ in formula 3.32 with $\Delta vh_{W}$ and then use that the integral of $h_{W}\omega_{0}^{n}$ vanishes, by Lemma 3.3, since the Futaki character of $X$ is assumed to vanish. It now follows that the function $t\mapsto\mathcal{J}(v\cdot\exp(-tJW))$ is strictly convex, precisely as in the proof of Prop 3.4, using that the measure $\omega_{v}\wedge\omega_{0}^{n-1}$ does not charge pluripolar subsets. Indeed, since $v\in\mathcal{E}^{1}(X,\omega)$ the measure defined as the product $\omega_{v}\wedge\omega_{0}^{n-1}$ of the current $\omega_{v}$ with the smooth form $\omega_{0}^{n-1}$ is identical to the non-pluripolar product of $\omega_{v}$ with $\omega_{0}^{n-1},$ which, by definition, does not charge pluripolar subsets (this identity follows, for example, from writing $v$ as a decreasing limit of elements in $\mathcal{E}^{1}(X,\omega)$ and using that both types of products are continuos wrt decreasing limits, since $v$ has full Monge-Ampère mass [25, Prop 2.1]). Next note that the pullback of $\mathcal{J}$ to $J\mathfrak{k}$ is proper and thus admits a critical point. Indeed, since $\mathcal{J}$ is comparable to $\mathcal{I}-\mathcal{J},$ which can be identified with the functional $E$ on $\mathcal{P}(X),$ the properness of $\mathcal{J}$ follows directly from the properness of $E,$ established in the course of the proof of Prop 3.4. Finally, when $\omega_{0}$ is Kähler–Einstein $\Delta h_{W}=\lambda_{1}h_{W}$ for any $W\in\mathfrak{k},$ where $\lambda_{1}$ denotes the smallest strictly positive eigenvalue of $\Delta$ [9, Thm 2.2]. Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.30, when $\omega_{0}$ is Kähler–Einstein. ∎
@@ 3.4. The reduced analytic stability thresholds of K-polystable log Fano curves
Given $t<\alpha(X,\Delta)_{0}$ it will be enough to show that the functional $\phi\mapsto\mathcal{I}_{t,\mu_{0}}(\phi_{\mu})$ is bounded from above on the subspace $K$ of $\text{PSH(}X,-K_{(X,\Delta)})$ consisting of all potentials $\phi$ of $\mu\in\mathcal{P}(X)_{0}.$ To this end first note that $K$ is a compact subspace of $\text{PSH(}X,-K_{(X,\Delta)}),$ endowed with the $L^{1}$–topology. Indeed, $\mathcal{P}(X)_{0}$ is a closed compact subset of the compact space $\mathcal{P}(X),$ hence compact and, as is well-known, the map $\mu\mapsto\phi_{\mu}$ is a homeomorphism. It will thus be enough to show that $\mathcal{I}_{t,\mu_{0}}$ is continuous on $K.$ But since $\mathcal{I}_{t+\delta,\mu_{0}}$ is pointwise finite on $K$ for $\delta$ a sufficiently small positive number, this is shown precisely as in the proof of [11, Cor A.2], using the corresponding local result of Demailly–Kollar.
@@
where $F(-\gamma)_{0}$ is defined by formula 3.17. Moreover, $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ and $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}$ when $w\neq 0.$
@@
Let us first show that for any log Fano curve $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$ We will adapt the argument in [11] (concerning the case $\mathcal{G}$ trivial) to the present setup. Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.18 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
@@
To this end recall that, $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ is independent of the reference $\phi_{0}$ (by Prop 3.6, using that the Futaki character of $(X,\Delta_{w})$ vanishes). Hence, we can take $\phi_{0}$ to be induced by the Fubini–Study metric. It will be enough to show that there exists a curve $\mu^{t}$ in $\mathcal{P}(X)_{0}$ such that for any $\gamma>2,$ $F_{-\gamma}(\mu^{t})\rightarrow$-$\infty$ as $t\rightarrow\infty.$ To this end first note that if $\mu$ is $S^{1}$–invariant under the standard $S^{1}$–action of $\mathbb{P}^{1}$ fixing 0 and $\infty$ and $\mathbb{Z}_{2}$–invariant under the $\mathbb{Z}_{2}$–action $(x_{1},x_{2},x_{3})\mapsto(-x_{1},-x_{2},-x_{3})$ then $\boldsymbol{m}(x_{1},x_{2},x_{3})=-\boldsymbol{m}(-x_{1},-x_{2},-x_{3})$ and it follows that $\mu\in\mathcal{P}(X)_{0}.$ We will construct $\mu^{t}$ as $\mu^{t}=dd^{c}\phi_{t}/V,$ where $\phi_{t}$ is an $S^{1}$–invariant geodesic in $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}},$ using the notation in [17] (recalled in the appendix), where such a curve $\phi_{t}$ is realized as the Legendre transform of a function $v_{t}(y)$ on the moment polytope P=[V/2,V/2]P=[-V/2,V/2] of $(X,-K_{(X,\Delta)})$ of the form
@@
where $v$ is a piecewise affine convex function on [1,1][-1,1] and $v_{0}$ is the Legendre transform of $\phi_{0}.$ It follows from well-known general results that such a curve $\phi_{t}$ is a geodesic ray in $\mathcal{H}(-K_{X})_{1,1}^{S^{1}}$ (see [52, Section 7]). Here we will take
@@
$F_{-\gamma}(\mu)=\gamma\frac{\mathcal{L}(v)}{V}+(1-\gamma)\frac{V}{2}\int_{P}\left|v^{\prime}\right|dy-\int_{P}v^{\prime\prime}dy+O(1),\,\,\,\frac{\mathcal{L}(v)}{V}\coloneqq\int_{\partial P}\frac{v}{2}-\int_{P}v\frac{dy}{V},$
where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 10.1, combined with 10.3). In particular,
@@
$F_{-\gamma}(\mu^{t})=\gamma t\frac{V}{4}+(1-\gamma)t\frac{V}{2}+O(1)=tV\left(\gamma\left(\frac{1}{4}-\frac{1}{2}\right)+\frac{1}{2}\right)+O(1),$
@@
The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 10.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
@@
The proof of the previous theorem shows that when $\gamma>\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ the functional $F_{-\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is destabilized by a geodesic ray, i.e. $F_{-\gamma}(\mu_{t})\rightarrow-\infty,$ as $t\rightarrow\infty,$ for a curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ of the form $\mu_{t}=\omega_{u_{t}}/V,$ where $u_{t}$ is a geodesic ray in $\mathcal{H}_{1,1}(X,\omega_{0}).$ This is reminiscent of the general existence/properness principle in [61, Section 3]. However, in the present setup the function $F_{-\gamma}(\omega_{u_{t}}/V)$ need not be convex in $t$ (since $\gamma>1)$ which is required in [61]. Moreover, our geodesics are required to satisfy the constraint $\int_{X}\boldsymbol{m}\omega_{u_{t}}=0.$